What is the probability of being murdered in the US?
Answer assumes the question refers to US residents and not immigrants, tourists, work permit bearers, etc.
According to the 2011 census, the US had a population 311,800,000 citizens.
According to the FBI website, there were 13,913 murders and non-negligent manslaughters in 2011.
Roughly speaking, in 2011 you had a 0.0045% chance of being murdered, if you were a US citizen in 2011.
judge
What is the probability of rolling 4 even numbers in a row on a 6 sided dice?
Even numbers: 2, 4, 6, the probability to get an ven number is 3/6 = 1/2
The probability to get 4 even number in a row is then (1/2)4 = 1/24 = 1 / 16
It is the same probability to get 4 odd numbers (1,3,5) in a row
One is a small sample size, but that's just my answer, you might want to ask more people.
What is the probability of picking a picture card at random from a standard deck of cards?
It is 3/13.
It is 11/13.
The empirical probability can only be determined by carrying out the experiment a very large number of times. Otherwise it would be the theoretical probability.
What is the standard deviation if all data points are different?
It depends on the data. The standard deviation takes account of each value, therefore it is necessary to know the values to find the sd.
How do you turn probability into a fraction?
An empirical estimate of the probability of an event is the ratio of the number of succesful outcomes to the total number of trials. By definition, the ratio is a fraction. However, there are many events for which the theoretical probability is related to irrational numbers. For example, it you randomly drop a pin on a floor of wooden boards, the probability that the pencil lies across a lateral join is related to pi. Being irrational, this cannot be expressed as a fraction.
How many simple random samples are possible of size 40 from a population of size 500?
Approximately 1.364*1060
Do you combine tails and heads?
No, because taken together, they account for all the possible outcomes and so combining them removes any informational value.
That is unless you believe that there is a significant probability that the coin will balance on its edge.
What does population distrubition mean?
In statistics, the population distribution of some variable within a population of some units is the set of values that the variable can take and how frequntly it takes that value.
Population need not have anything to do with people. It is simply a way of referring to the totality of the things that you are studying. For example, if you were studying the number of apples on each tree in an apple orchard, your population would be the apple trees in the orchard, the variable would be the number of apples per tree and the population distribution would be the number of trees that had 0 apples, the number that had 1 apple, 2 apples and so on. To make that information more manageable you may collapse the distribution to 0 apples, 1 to 25 apples, 26 to 50 apples and so on.
What is the probability of choosing a card that is not a ace but that is divisible by 2?
It is 5/13 if you assume that Q is not divisible by 2.
What is the number of times an event or value occurs in a given period?
The frequency of the event over that time period.
There is no such thing as "gentic makeup". Perhaps you mean genetic, instead. There is a probabilistic element in genetic make up, but in some cases the probabilistic bit is overridden by recessive genes.
What is the probability you roll a die and get a 2 and an even number?
The probability of getting a 2 is (1/6) as there are 6 sides, the probability of then getting an even number 2,4, or 6 is 3/6 or (1/2). Therefore those 2 even occurring together are 1/6 X 1/2 =1/12 or 8.5%.
What is the most probable ion formed from AI?
AI is not an element or compound and so cannot form an ion.
What are the odds of having a baby with 3 very rare diseases?
If you know nothing about the parents' risk for these diseases or other contributing factors (race, environment, etc.), then the answer depends solely on the probabilities of getting any one of those diseases.
Let's say the risks are P1, P2 and P3 where P1 is the probability of the baby having rare disease #1 and so on. A probability of 1 is a sure thing and 0 means no chance, so the probability of a rare disease is a number slightly greater than 0.
The risk of getting disease #1 AND #2 AND #3 is the product: P1xP2xP3.
If, for example, the each risk is one in a million (P1=P2=P3=0.000001), then the odds of having a baby with all three is 1 in 1,000,000,000,000,000,000.
The math changes severely if one of the parents is at risk for the disease. Let's say the mother has one gene that causes the disease #1. Then P1 is no longer 1 in a million, it is 0.5. If the parent's racial group is afflicted with P2=0.001 as well, then P=0.5x0.001x0.000001=0.000000005, which is 500,000 the risk of others.
-1.43 (approx)
The ideal sample size depends on a number of factors:
What is the best deck in elements the card game?
The Deck that i run is really basic, all you need is about 500 coins to make it, im going to stop blabbing and give you the deck build below!
12 obsidian pillars
6 devour
6 minor Vampire
3 Darkness Dragons (Optional, it slows the deck down SLIGHTLY, but increases your attack power)
6 Aether Pillars
4 Fractal
2 nightfall
2 Steal
1 Dimensional Shield
There you have it! With the Dragons in the deck, you have a very basic but effective DDD Deck! (Darkness Dragon Destruction) or if you choose to have them out of the deck its a DD, Devour Drainer deck. Based on your strategy on this deck, it can be either
Darkness Dragon Destruction
Devour Drain
Vampire Swarm
Because of the Fractal, you can fill your hand with any card on the field that you want, fill the field with dragons, Vampires, or Devour's depending on your opponent! Enjoy!!
When a six-sided die is thrown what is the probability of getting a 1?
There is a one out of six chance.
What is the probability that you will choose a club or a jack in a standard deck of 52 cards?
It is 16/52 = 4/13
How many ways can the word hospital be arranged?
Eight letters, so....
8!
= 40320 different ways
=================